# Adaptive Linearly Constrained Minimum Variance Volumetric Active Noise Control

### v1

@@ Abstract
Index Terms:
Active Noise Control, Beamforming, Optimization, Adaptive Filters, Sensor Array Processing
@@ I Introduction
Volumetric noise control aims to reduce unwanted sound energy throughout a three-dimensional space [3], rather than at a limited number of microphone locations. Unlike conventional ANC approaches that minimize noise at specific spatial points (e.g., near the ears or at sensor positions), volumetric control seeks to attenuate noise over an extended region—such as an entire cabin, room, or workspace. This goal typically requires a spatially distributed arrangement of microphones and speakers. Prior methods have attempted to achieve this through multi-point FxLMS control, kernel-interpolation techniques to estimate the acoustic field within a volume [4], or virtual sensing strategies [5].
@@
In array signal processing [6], constrained optimization techniques [7] have proven highly effective, especially for beamforming tasks. The Linearly Constrained Minimum Variance (LCMV) beamformer, in particular, minimizes the output power (variance) of an array while satisfying linear constraints on its response [6, 8]. These constraints are typically designed to preserve signals from desired directions or suppress interference from known sources. Although LCMV methods are primarily used in spatial filtering with sensor arrays, the core idea—minimizing variance subject to linear constraints—can be adapted to design temporal adaptive filters for ANC.
This paper proposes a novel time-domain adaptive algorithm for volumetric ANC, framed within the beamforming paradigm [9]. A frequency-domain LCMV formulation for ANC was introduced by Huang et al. [10], who derived an ideal solution under the assumption of known primary and secondary paths using acausal, infinite-length filters. However, the assumption of a known primary path is often impractical. The method presented in this work eliminates this assumption and learns the necessary information directly from the data.
A key advantage of the proposed algorithm is its operation entirely in the time domain, which is beneficial in latency-sensitive applications where frequency-domain methods may introduce delay. The approach utilizes the Frost Generalized Sidelobe Canceller (GSC) structure [11, 12] to implement the adaptive LCMV filter. The GSC architecture decomposes the constrained optimization into two paths: a fixed path that enforces the constraints and an adaptive path that operates on components orthogonal to the constraint space. This structure enables the use of standard unconstrained LMS-type updates in the adaptive path, simplifying computation while implicitly satisfying the LCMV constraints.
By combining the LCMV optimization objective with the low-latency, time-domain GSC structure, the proposed algorithm provides effective volumetric noise control with explicitly enforced linear constraints. Furthermore, the method naturally extends to kernel-interpolated fields. This provides a flexible design trade-off: Certain spatial locations—such as those near a listener’s ears—can be strictly constrained to achieve reliable noise attenuation, while the rest of the volume can be optimized under a minimum variance objective. This hybrid approach ensures strong performance at critical points without sacrificing global noise reduction. Simulation and experimental results demonstrate the effectiveness of the proposed adaptive LCMV-ANC approach.
@@ IV-B Adaptive Algorithm
In a real-time active noise control scenario, one must update the filter coefficients online to adapt to the time-varying noise. In this section, we derive an adaptive algorithm based on Frost’s generalized sidelobe canceler (Frost-GSC). In particular, we apply stochastic gradient descent (SGD) to the Lagrangian in the optimization problem. Explicitly, the update rule may be expressed as,
@@ V-B Experimental Results
Figure 4 shows the convergence curves for the proposed method, two-point FxLMS and multi-point FxLMs. The proposed method is able to reduce noise at the constrained location more effectively than multi-point FxLMS but is unable to match two-point FxLMS. Figure 5 shows the power spectral density (PSD), plotted in decibels (dB), at the two in-ear microphones for each algorithm. The PSD is limited to the 0–16 kHz frequency range for clarity. The proposed method clearly achieves better satisfaction of the linear constraints compared to the multi-point FxLMS, with noticeably improved performance across a broad frequency range. Although the proposed method trades off some performance at secondary control locations due to the inclusion of a minimum variance component, it still achieves results within 1 dB of the multi-point FxLMS at those locations. Figure 6 shows the average convergence behavior and the average PSD (also in dB, limited to 0–16 kHz) at the secondary control microphones. These results demonstrate that the proposed method delivers superior performance at the primary control locations, while maintaining performance at the secondary locations comparable to that of the multi-point FxLMS.
@@ VI Conclusion
The time-domain formulation and adaptive algorithm presented in this work reframe active noise control as a problem of informed spatial trade-offs rather than uniform suppression. By embedding linear constraints directly into the optimization process via a Lagrangian approach, our method allows precise control over where in space noise reduction is prioritized, reflecting a shift from purely reactive to strategically targeted spatial filter design. The demonstrated gains at constrained locations without degrading overall performance affirm the value of this perspective. This framework allows for effective spatial control, particularly when coupled with future extensions such as kernel-based noise field interpolation, which promises to further reduce the burden of spatial sampling while enhancing control fidelity.

### v2

# Adaptive Linearly Constrained Minimum Variance Framework for Volumetric Active Noise Control

@@ I Introduction
Volumetric noise control aims to reduce unwanted sound energy throughout a three-dimensional space [3, 4], rather than at a limited number of microphone locations. Unlike conventional ANC approaches that minimize noise at specific spatial points (e.g., near the ears or at sensor positions), volumetric control seeks to attenuate noise over an extended region—such as an entire cabin, room, or workspace. This goal typically requires a spatially distributed arrangement of microphones and speakers. Prior methods have attempted to achieve this through multi-point FxLMS control, kernel-interpolation techniques to estimate the acoustic field within a volume [5], or virtual sensing strategies [6, 7].
@@
In array signal processing [8], constrained optimization techniques [9] have proven highly effective, especially for beamforming tasks. The Linearly Constrained Minimum Variance (LCMV) beamformer, in particular, minimizes the output power (variance) of an array while satisfying linear constraints on its response [8, 10]. These constraints are typically designed to preserve signals from desired directions or suppress interference from known sources. Although LCMV methods are primarily used in spatial filtering with sensor arrays, the core idea—minimizing variance subject to linear constraints—can be adapted to design temporal adaptive filters for ANC. This paper proposes a novel time-domain adaptive algorithm for volumetric ANC, framed within the beamforming paradigm [11]. A frequency-domain LCMV formulation for ANC was introduced by Huang et al. [12], who derived an ideal solution under the assumption of known primary and secondary paths using acausal, infinite-length filters. However, the assumption of a known primary path is often impractical. The method presented in this work eliminates this assumption and learns the necessary information directly from the data.
A key advantage of the proposed algorithm is its operation entirely in the time domain [13], which is beneficial in latency-sensitive applications where frequency-domain methods may introduce delay. By combining the LCMV optimization objective with the low-latency, time-domain generalized sidelobe canceler structure, the proposed algorithm provides effective volumetric noise control with explicitly enforced linear constraints. Furthermore, the method naturally extends to kernel-interpolated fields. This provides a flexible design trade-off: Certain spatial locations—such as those near a listener’s ears—can be strictly constrained to achieve reliable noise attenuation, while the rest of the volume can be optimized under a minimum variance objective. In contrast to prior works—such as the dual zone formulation [14], the directional ANC [15, 16] with minimax [17], and spatially selective systems to preserve certain signals [18, 19, 20, 21, 22], a fundamentally different problem, we pose the problem for controlled filter design in a volume which may be coupled with directional or spatially selective constraints and extends to an arbitrary number of zones.
This hybrid approach ensures strong performance at critical points without sacrificing global noise reduction. Simulation and experimental results demonstrate the effectiveness of the proposed adaptive LCMV-ANC approach.
@@ IV-B Adaptive Algorithm
In a real-time active noise control scenario, one must update the filter coefficients online to adapt to the time-varying noise. We apply stochastic gradient descent (SGD) to the Lagrangian in the optimization problem. Explicitly, the update rule may be expressed as,
@@ V-B Experimental Results
Figure 4 shows the convergence curves for the proposed method, two-point FxLMS and multi-point FxLMs. The proposed method is able to reduce noise at the constrained location more effectively than multi-point FxLMS but is unable to match two-point FxLMS. Figure 5 shows the power spectral density (PSD), plotted in decibels (dB), at the two in-ear microphones for each algorithm. The PSD is limited to the 0–16 kHz frequency range for clarity. The proposed method clearly achieves better satisfaction of the linear constraints compared to the multi-point FxLMS, with noticeably improved performance across a broad frequency range. The reduced low frequency performance is because the acoustic head simulator’s HRTF must be estimated along with the room transfer function for those points causing a significant change in the acoustic manifold in that region. Although the proposed method trades off some performance at secondary control locations due to the inclusion of a minimum variance component, it still achieves results within 1 dB of the multi-point FxLMS at those locations. Figure 6 shows the average convergence behavior and the average PSD (also in dB, limited to 0–16 kHz) at the secondary control microphones. These results demonstrate that the proposed method delivers superior performance at the primary control locations, while maintaining performance at the secondary locations comparable to that of the multi-point FxLMS.
@@ VI Conclusion
The time-domain formulation and adaptive algorithm presented in this work reframe active noise control as a problem of informed spatial trade-offs rather than uniform suppression. By embedding linear constraints directly into the optimization process via a Lagrangian approach, our method allows precise control over where in space noise reduction is prioritized, reflecting a shift from purely reactive to strategically targeted spatial filter design. The demonstrated gains at constrained locations without degrading overall performance affirm the value of this perspective. This framework of LCMV active noise control allows system designers flexibility in how degrees of freedom are allocated for the control of the acoustic field. In particular, our method readily extends to any spatial filtering scheme for active noise control and allows for effective spatial control, particularly when coupled with future extensions such as kernel-based noise field interpolation, which promises to further reduce the burden of spatial sampling while enhancing control fidelity.